OpenStudy (anonymous):

Josh and his friends bought vanilla wafers for $4 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $45 to buy a total of 27 packets of wafers of the two varieties.

OpenStudy (anonymous):

Part A: Write a system of equations that can be solved to find the number of packets of vanilla wafers and the number of packets of chocolate wafers that Josh and his friends bought at the carnival. Define the variables used in the equations. (5 points) Part B: How many packets of chocolate wafers and vanilla wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer. (5 points)

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

will fan u

OpenStudy (texaschic101):

v + c = 27 --- there are 27 vanilla and choco packs of wafers 4v + 1c = 45 -- $4 for each v bought + $1 for each c bought = $45 these are your 2 equations that you will use....do you need help in solving these now ?

OpenStudy (texaschic101):

v + c = 27 v = 27 - c now sub in 27 - c in for v back into the other equation 4v + 1c = 45 4(27 - c) + 1c = 45 --- so basically, you distribute through the parenthesis and then solve for c. Once c is found, sub that value back into v + c = 27 to find v

OpenStudy (anonymous):

how to u do the second one